OpenStudy (idku):

I got a question about vector problem. (Please, I just started don't through hard stuff at me).

OpenStudy (idku):

Two balls are thrown from the same height as shown. The \(\color{red}{\rm red}\) ball is thrown with velocity \(\color{red}{ \vec v}\) in the vertical direction. The \(\color{blue}{\rm blue}\) ball is thrown with twice the speed of the red ball at an angle of 30\(^\circ\) with respect to the horizontal. Which ball will go higher? 

OpenStudy (idku):

The choices they gave me are: A) the red ball B) the blue ball C) the maximum height for both balls is the same D) it depends on the value of v but I really want to figure how to do these.

OpenStudy (michele_laino):

hint: the vertical speed of the red ball is V whereas the vertical speed of the blue ball is 2V*sin(30)=2*V/2=V

OpenStudy (idku):

V \(?\) 2V \(\sin(30)\) V \(?\) 2V \(\sin(30)\) V \(?\) 2V \((1/2)\) V \(=\) V \(\)

OpenStudy (idku):

So they will reach the same height, right?

OpenStudy (michele_laino):

yes!

OpenStudy (idku):

(regardless of the V)

OpenStudy (michele_laino):

yes!

OpenStudy (idku):

I mean V is still greater than zero, just from the problem. Not that we even need to know this. Thank you!! Just one more thing, so for a horizontal motion of some vector \(\bf v\), with a degree of \(\theta\), the ball will go a horizontal distance D of: \({\rm D}={\bf v}\times \sin(\theta)\)

OpenStudy (idku):

(right? )

OpenStudy (michele_laino):

in order to compute the horizontal distance, we have to consider the horizontal component of the speed of the blue ball, whose magnitude is: 2*V cos(30)=sqrt(3)*V

OpenStudy (michele_laino):

\[\Large {v_x} = 2V\cos 30 = 2V \cdot \frac{{\sqrt 3 }}{2} = \sqrt 3 V\]

OpenStudy (idku):

So, in general, if we had:  then the horizontal distance is: \({\rm D}={\bf v}\times \sin(\theta) \times {\bf v} \cos(\theta)\)

OpenStudy (michele_laino):

no, the horizontal distance is: \[\Large D = {v_x}\Delta t\] where \Delta t is the total flying time

OpenStudy (idku):

and Vx is what exactly?

OpenStudy (michele_laino):

Vx is: \[\Large {v_x} = V\cos \theta \]

OpenStudy (idku):

\(\large {\rm D_{horizontal }}=\left(\Delta t\right) {\bf v} \cos(\theta)\)